# A systematically-selected sample of luminous, long-duration, ambiguous nuclear transients

### v1

@@ Abstract
We present a search for luminous, long-duration ambiguous nuclear transients (ANTs) similar to the unprecedented discovery of the extreme, ambiguous event AT2021lwx with a d rise time and luminosity erg s-1. We use the Lasair transient broker to search Zwicky Transient Facility (ZTF) data for transients lasting more than one year and exhibiting smooth declines. Our search returns 59 events, seven of which we classify as ANTs assumed to be driven by accretion onto supermassive black holes. We propose the remaining 52 are stochastic variability from regular supermassive black hole accretion rather than distinct transients. We supplement the seven ANTs with three nuclear transients in ZTF that fail the light curve selection but have clear single flares and spectra that do not resemble typical AGN. All but one of these 10 ANTs have a mid-infrared flare from an assumed dust echo, implying the ubiquity of dust around the black holes giving rise to ANTs. No events are more luminous than AT2021lwx, but one (ZTF19aamrjar) has twice the duration and a higher integrated energy release. On the other extreme, ZTF20abodaps reaches a luminosity close to AT2021lwx with a rise time d and that fades smoothly in d. We define a portion of rise-time versus flare amplitude space that selects ANTs with per cent purity against variable active galactic nuclei. We calculate a volumetric rate of Mpc-1 yr-1, consistent with the events being caused by tidal disruptions of intermediate and high-mass stars.
@@
AGN are inherently variable across the entire electromagnetic spectrum: in the ultra-violet–optical–near-infrared (UVOIR) the variability is stochastic on timescales from seconds to years, and is typically limited to amplitudes mag on these timescales (e.g. Vanden Berk et al., 2004; MacLeod et al., 2010; Caplar et al., 2017; Sheng et al., 2022). Their variations are often described by a stochastic process called a damped random walk (DRW; Kelly et al. 2009). The presence of short-term variability in the optical is challenging to the canonical thin, viscous accretion disk model (e.g. Lawrence, 2018; Antonucci, 2018) and indicates that the disk reprocesses higher frequency (i.e. X-ray) emission from a small central region (e.g. Clavel et al., 1992; McHardy et al., 2016). Such reprocessing, however, fails to explain the increasing number of AGN showing large-amplitude changes on short timescales: it is estimated that up to 30-50 per cent of quasars show variability at the mag level over baselines of 15 yr (Rumbaugh et al., 2018), dubbed ‘extreme variability quasars’ (EVQs). Another growing class of AGN is that showing distinct spectral changes on years – decade timescales: the (dis)appearance of broad emission lines (e.g. LaMassa et al., 2015; MacLeod et al., 2016), sometimes temporally coincident with a change in flux. So-called ‘changing-look AGN’ (CLAGNs) challenge theories of AGN emission mechanisms even more (for a review see Ricci & Trakhtenbrot 2022). Correlations between the continuum luminosity and broad emission line appearance indicate that changes in the accretion flow drive CLAGNs, also called changing state AGNs111Note that changing-look has also been used to describe those AGN which change X-ray properties, which can also be denoted changing obscuration AGN (Ricci & Trakhtenbrot, 2022). In one extreme case, a non-variable AGN appears to ‘switch on’ its variability over the course of a few years (Ridley et al., 2023) while another changes its spectral state on a timescale of months (Trakhtenbrot et al., 2019b).
@@
The distinction between ANTs and AGN flares, CLAGNs, and TDEs, is not clearly defined. Indeed, TDEs are not excluded from occurring in AGN. ANT light curves appear to fall into two categories. Some ANT light curves are very smoothly evolving, rising slowly and decline monotonically even slower, including AT2019brs (Frederick et al., 2021), Gaia16aaw/AT2016dbs and Gaia18cdj/AT2018fbb (Hinkle et al., 2024), and the most energetic transient ever discovered, AT2021lwx (Wiseman et al., 2023; Subrayan et al., 2023). This behaviour appears to be a slower version of the evolution of most TDEs. Other ANTs, although displaying a single overall flare, show variability superimposed over their long-term light curves, such as AT2019fdr (Frederick et al., 2021; Reusch et al., 2022; Pitik et al., 2022), ASASSN-15lh (Leloudas et al., 2016)222Interpreted in Leloudas et al. (2016) and Krühler et al. (2017) as a TDE, but by Dong et al. (2016) as a SLSN., ASASSN-17jz (Gromadzki et al., 2019), ASASSN-18jd (Neustadt et al., 2020), PS16dtm (e.g. Blanchard et al., 2017; Petrushevska et al., 2023), AT2017bgt (Trakhtenbrot et al., 2019a) and Swift J2219510-484240 (Oates et al., 2024). While their light curves are similar to TDEs, ANTs are spectroscopically more similar to AGN with high equivalent width emission lines (particularly from the hydrogen Balmer series), while TDE emission lines are much broader but weaker. A subset of ANTs show features associated with Bowen fluorescence such as He ii and N iii, indicative of a steep far-UV or soft X-ray source such as accretion disk, and are termed Bowen fluorescence flares (BFFs; Trakhtenbrot et al. 2019a). Some events classified as TDEs also have signatures of Bowen fluorescence, complicating the distinction between TDEs and ANTs further but hinting that the emission mechanisms are similar, irrespective of the source of accretion (Leloudas et al., 2019; Blagorodnova et al., 2017; van Velzen et al., 2021). Frederick et al. (2021) presented a small sample of ANTs in known narrow-line Seyfert 1 (NLSy1) AGN, a class of Seyfert galaxies accreting close to the Eddington limit. PS16dtm also occurred in a NLSy1 (Blanchard et al., 2017), while PS1-10adi also showed similar features (Kankare et al., 2017), raising the possibility that ANTs preferentially occur in such systems.
@@
An extra piece of the TDE - ANT - AGN puzzle is provided by mid-infrared (MIR) observations that probe hot dust. MIR flares, which tend to lag the UVOIR, accompany many of the UVOIR-discovered ANTs (Hinkle, 2022; Petrushevska et al., 2023; Wiseman et al., 2023; Oates et al., 2024), as well as existing alone without optical counterparts (Jiang et al., 2021a; Wang et al., 2022). MIR flares are observed less frequently in TDEs (van Velzen et al., 2016), although the population of MIR flares without optical counterparts (Jiang et al., 2021a; Masterson et al., 2024) could be caused by fully obscured TDEs. The interpretation is that these MIR-loud ANTs occur in nuclei with circumnuclear dust, typically described as the AGN ‘torus’. Nevertheless, some ANTs show no sign of previous AGN activity and optical spectra distinct from typical AGN (e.g. Oates et al., 2024).
@@
In Section 2 we introduce the search for ANTs in the Zwicky Transient Facility data stream. In Section 3 we present the individual ANTs and their shared properties. In Section 4 we compare the variability properties of ANTs and AGN. We discuss the implications on the nature of ANTs in Section 5, and conclude in Section 6. Where relevant we assume a spatially flat -cold dark matter () cosmology with km s-1 Mpc-1, and . Unless otherwise stated, uncertainties are presented at the level.
Cut | Number cut | Number remaining |
@@
Cut | Number cut | Number remaining |
@@
ZTF ID | RA | Dec | IAU Name | Crossmatches | Host -mag | Redshift | Light curve classification |
ZTF20abrbeie | 318.45173 | 27.43066 | AT2021lwx | ATLAS20bkdj, PS22iin | No detected host | 0.9945 a | ANT |
ZTF19aamrjar | 272.33963 | 25.30919 | - | ATLAS19mmu | 18.3 | 0.697 b | ANT |
ZTF20abodaps | 359.80815 | -17.69658 | AT2020afep | ATLAS20vrw | 19.8 | 0.607 b | ANT |
ZTF18aczpgwm | 33.79096 | 7.16453 | AT2019kn | ATLAS19bdfo, Gaia19abv | 19.1 | 0.4279 b | ANT |
ZTF21abxowzx | 324.33304 | -10.753312 | AT2021yzu | ATLAS21bjoi | 20.3 | 0.419 c | ANT |
ZTF20aaqtncr | 251.36440 | 6.29627 | AT2021fez | ATLAS20pzv, Gaia21bgs | 19.7 | 0.368 b | Unclear |
ZTF19aailpwl | 216.94340 | 29.51061 | AT2019brs | ATLAS19fyh, Gaia19axp | 19.3 | 0.3736 d | ANT |
ZTF20acvfraq | 349.72403 | -10.58489 | AT2020adpi | ATLAS20bjzp, Gaia21aid | 19.7 | 0.26 e | ANT |
ZTF19aatubsj | 257.27857 | 26.85569 | AT2019fdr | ATLAS19lkd | 18 | 0.2666 f,d | ANT |
ZTF20aanxcpf | 15.16508 | 39.70842 | AT2021loi | ATLAS21qje | 18 | 0.083 g | ANT |
@@ 2.1.1 Nuclear transients
ANTs are, by definition, located in galaxy nuclei. To select nuclear transients we use the sherlock sky context software (Smith et al., 2020) running in Lasair. The ZTF search area is fully covered by the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS1) science consortium surveys (Chambers et al., 2016; Wainscoat et al., 2016): sherlock uses the Pan-STARRS catalogues and probabilistic classifications of unresolved point sources (Tachibana & Miller, 2018) to determine a host galaxy and the transient location within that host. We select all objects that are predicted as Nuclear Transients (i.e. transients consistent with a galactic nucleus) or AGN (transients consistent with a galactic nucleus that are listed in an AGN catalogue). Note that most transient searches exclude known AGN, but we allow them to avoid excluding bona-fide transient events associated with actively accreting SMBHs. We also include all objects predicted as SNe (transients not entirely consistent with a galaxy nucleus) with an angular separation of from the catalogued centre of a galaxy. Finally, we include “Orphan" transients, which have no associated host galaxy in deeper survey imaging (as was the case for AT2021lwx). Including transient that appear hostless prevents the systematic exclusion of events at high-redshift, extremely dusty, or uncatalogued dwarf galaxies.
@@ 2.2 Photometric selection of ANTs
Duration: ANTs are long-lived transients, and we require that the most recent detection in at least one filter must be at least 1 yr after the first detection.
@@
Linear decline: Any remaining AGN and variable stars should show stochastic variability on timescales shorter than the typical decline of an ANT. Single transient events on the other hand show power-law or exponential decays, which are linear in magnitude space. We fit the declines with a linear model and quantify the goodness-of-fit to the linear model with the Pearson parameter. We set an arbitrary threshold of for a decline to be accepted as linear.
After all the nominal selection criteria are applied, we are left with 58 nuclear light curves and one orphan (ZTF20abrbeie=AT2021lwx), which we denote the ‘gold selection’. Of the ANTs published by Frederick et al. (2021) we recover ZTF19aailpwl=AT2019brs (the other three are too ‘bumpy’). To explore ANTs with different light curve shapes, we subsequently trial a set of different selection criteria, which are described in Appendix A. Briefly: we conduct a search for faster-declining events by relaxing the maximum duration to 180 d (the ‘faster’ sample); for less smooth events by relaxing the linear decline correlation coefficient to (the ‘bumpier’ sample); and the decline time to be longer than the rise time (the ‘platinum’ sample). The resulting samples and their ZTF light curve properties are available online555https://github.com/wisemanp/ANTs-Nest. For the 59 gold events, we obtain forced photometry for the full ZTF survey duration via the ZTF forced photometry service (Masci et al., 2019) for visual inspection.
@@ 2.2.1 Visual inspection
We visually inspect all 59 lightcurves and assign visual classifications of ‘ANT’, ‘AGN’, or ‘unclear’ according to the flowchart in Fig 2. These classifications are predominantly based on pre-and post-peak variability. In particular, the forced photometry allows us to view any variability that occurred before the candidate passed the alert threshold in ZTF. Eight events were visually classified as ANTs, and one as Uncertain.
2.2.2 Spectroscopic follow-up
Of the 59 ‘gold’ light curves, 17 were previously reported as transients to the Transient Name Server (TNS)666https://www.wis-tns.org/, and a handful have been publicly spectroscopically classified. Note that ANT is not a spectroscopic classification. Two ZTF identifiers correspond to the same transient (ZTF22aafujzv=ZTF22aadesap=AT2022fpx) which is spectroscopically classified as a TDE at . However, the spectra of AT2022fpx do not resemble the main classes of TDE defined by Arcavi et al. (2014); van Velzen et al. (2021) and the object has recently been identified as an extreme coronal line emitter (Koljonen et al., 2024) ZTF22aanwibf = SN2022mma is a type IIn SN at , and on visual inspection appears in a star-forming region of a spiral arm of its host that is catalogued as a separate source in PS1 and hence passes the s̈election. The remaining ANT candidate host environments were visually inspected in the lasair broker to verify they are truly nuclear.
Of the eight ANTs and one Uncertain event passing visual inspection, three have existing public classifications: ZTF19aailpwl=AT2019brs is an ANT (Frederick et al., 2021), and ZTF21abxowzx=AT2021yzu is spectroscopically classified as an AGN at : in Section 3.1 we classify this as an ANT. ZTF20abrbeie=AT2021lwx (Wiseman et al., 2023; Subrayan et al., 2023) passes our selection by construction. Further to public classifications, we obtained spectra for four events that passed visual inspection: ZTF18aczpgwm=AT2019kn, ZTF19aamrjar, ZTF20abodaps=AT2020afep, ZTF20aaqtncr=AT2021fez. These spectra are described in detail in Section 3.1.

### v2

@@ Abstract
We present a search for luminous, long-duration ambiguous nuclear transients (ANTs) similar to the unprecedented discovery of the extreme, ambiguous event AT2021lwx with a d rise time and luminosity erg s-1. We use the Lasair transient broker to search Zwicky Transient Facility (ZTF) data for transients lasting more than one year and exhibiting smooth declines. Our search returns 59 events, seven of which we classify as ANTs assumed to be driven by accretion onto supermassive black holes. We propose the remaining 52 are stochastic variability from regular supermassive black hole accretion rather than distinct transients. We supplement the seven ANTs with three nuclear transients in ZTF that fail the light curve selection but have clear single flares and spectra that do not resemble typical AGN. All of these 11 ANTs have a mid-infrared flare from an assumed dust echo, implying the ubiquity of dust around the black holes giving rise to ANTs. No events are more luminous than AT2021lwx, but one (ZTF19aamrjar) has twice the duration and a higher integrated energy release. On the other extreme, ZTF20abodaps reaches a luminosity close to AT2021lwx with a rise time d and that fades smoothly in d. We define a portion of rise-time versus flare amplitude space that selects ANTs with per cent purity against variable active galactic nuclei. We calculate a volumetric rate of Mpc-1 yr-1, consistent with the events being caused by tidal disruptions of intermediate and high-mass stars.
@@
AGN are inherently variable across the entire electromagnetic spectrum: in the ultra-violet–optical–near-infrared (UVOIR) the variability is stochastic on timescales from seconds to years, and is typically limited to amplitudes mag on these timescales (e.g. Vanden Berk et al., 2004; MacLeod et al., 2010; Caplar et al., 2017; Sheng et al., 2022). Their variations are often described by a stochastic process called a damped random walk (DRW; Kelly et al. 2009). The presence of short-term variability in the optical is challenging to the canonical thin, viscous accretion disk model (e.g. Lawrence, 2018; Antonucci, 2018) and indicates that the disk reprocesses higher frequency (i.e. X-ray) emission from a small central region (e.g. Clavel et al., 1992; McHardy et al., 2016). Such reprocessing, however, fails to explain the increasing number of AGN showing large-amplitude changes on short timescales: it is estimated that up to 30-50 per cent of quasars show variability at the mag level over baselines of 15 yr (Rumbaugh et al., 2018), dubbed ‘extreme variability quasars’ (EVQs). Another growing class of AGN is that showing distinct spectral changes on years – decade timescales: the (dis)appearance of broad emission lines (e.g. LaMassa et al., 2015; MacLeod et al., 2016), sometimes temporally coincident with a change in flux. So-called ‘changing-look AGN’ (CLAGNs) challenge theories of AGN emission mechanisms even more (for a review see Ricci & Trakhtenbrot 2023). Correlations between the continuum luminosity and broad emission line appearance indicate that changes in the accretion flow drive CLAGNs, also called changing state AGNs111Note that changing-look has also been used to describe those AGN which change X-ray properties, which can also be denoted changing obscuration AGN (Ricci & Trakhtenbrot, 2023). In one extreme case, a non-variable AGN appears to ‘switch on’ its variability over the course of a few years (Ridley et al., 2024) while another changes its spectral state on a timescale of months (Trakhtenbrot et al., 2019b).
@@
The distinction between ANTs and AGN flares, CLAGNs, and TDEs, is not clearly defined. Indeed, TDEs are not excluded from occurring in AGN. ANT light curves appear to fall into two categories. Some ANT light curves are very smoothly evolving, rising slowly and decline monotonically even slower, including AT2019brs (Frederick et al., 2021), Gaia16aaw/AT2016dbs and Gaia18cdj/AT2018fbb (Hinkle et al., 2024), and the most energetic transient ever discovered, AT2021lwx (Wiseman et al., 2023; Subrayan et al., 2023). This behaviour appears to be a slower version of the evolution of most TDEs. Other ANTs, although displaying a single overall flare, show variability superimposed over their long-term light curves, such as AT2019fdr (Frederick et al., 2021; Reusch et al., 2022; Pitik et al., 2022), ASASSN-15lh (Leloudas et al., 2016)222Interpreted in Leloudas et al. (2016) and Krühler et al. (2017) as a TDE, but by Dong et al. (2016) as a SLSN., ASASSN-17jz (Holoien et al., 2022), ASASSN-18jd (Neustadt et al., 2020), PS16dtm (e.g. Blanchard et al., 2017; Petrushevska et al., 2023), AT2017bgt (Trakhtenbrot et al., 2019a) and Swift J2219510-484240 (Oates et al., 2024). While their light curves are similar to TDEs, ANTs are spectroscopically more similar to AGN with high equivalent width emission lines (particularly from the hydrogen Balmer series), while TDE emission lines are much broader but weaker. A subset of ANTs show features associated with Bowen fluorescence such as He ii and N iii, indicative of a steep far-UV or soft X-ray source such as accretion disk, and are termed Bowen fluorescence flares (BFFs; Trakhtenbrot et al. 2019a). Some events classified as TDEs also have signatures of Bowen fluorescence, complicating the distinction between TDEs and ANTs further but hinting that the emission mechanisms are similar, irrespective of the source of accretion (Leloudas et al., 2019; Blagorodnova et al., 2017; van Velzen et al., 2021). Frederick et al. (2021) presented a small sample of ANTs in known narrow-line Seyfert 1 (NLSy1) AGN, a class of Seyfert galaxies accreting close to the Eddington limit. PS16dtm also occurred in a NLSy1 (Blanchard et al., 2017), while PS1-10adi also showed similar features (Kankare et al., 2017), raising the possibility that ANTs preferentially occur in such systems.
@@
An extra piece of the TDE - ANT - AGN puzzle is provided by mid-infrared (MIR) observations that probe hot dust. MIR flares, which tend to lag the UVOIR, accompany many of the UVOIR-discovered ANTs (Hinkle, 2024; Petrushevska et al., 2023; Wiseman et al., 2023; Oates et al., 2024), as well as existing alone without optical counterparts (Jiang et al., 2021a; Wang et al., 2022). MIR flares are observed less frequently in TDEs (van Velzen et al., 2016), although the population of MIR flares without optical counterparts (Jiang et al., 2021a; Masterson et al., 2024) could be caused by fully obscured TDEs. The interpretation is that these MIR-loud ANTs occur in nuclei with circumnuclear dust, typically described as the AGN ‘torus’. Nevertheless, some ANTs show no sign of previous AGN activity and optical spectra distinct from typical AGN (e.g. Oates et al., 2024).
@@
In Section 2 we introduce the search for ANTs in the Zwicky Transient Facility data stream. In Section 3 we present the individual ANTs and their shared properties. In Section 4 we compare the variability properties of ANTs and AGN. We discuss the implications on the nature of ANTs in Section 5, and conclude in Section 6. Where relevant we assume a spatially flat -cold dark matter () cosmology with km s-1 Mpc-1, and . Unless otherwise stated, uncertainties are presented at the level. Magnitudes are presented in the AB system (Oke & Gunn, 1983).
Selection | Number cut | Number remaining |
@@
Visual inspection | 51 | 7 |
Selection | Number cut | Number remaining |
@@
Visual inspection | 0 | 1 |
ZTF ID | RA | Dec. | IAU name | Crossmatches | Host -mag | Redshift |
Photometric ANTs | ||||||
ZTF20abrbeie | 318.45173 | 27.43066 | AT2021lwx | ATLAS20bkdj, PS22iin | No detected host | 0.9945 abc |
ZTF19aamrjar | 272.33963 | 25.30919 | - | ATLAS19mmu | 18.3 | 0.697 d |
ZTF20abodaps | 359.80815 | -17.69658 | AT2020afep | ATLAS20vrw | 19.8 | 0.607 d |
ZTF18aczpgwm | 33.79096 | 7.16453 | AT2019kn | ATLAS19bdfo, Gaia19abv | 19.1 | 0.4279 d |
ZTF21abxowzx | 324.33304 | -10.753312 | AT2021yzu | ATLAS21bjoi | 20.3 | 0.419 e |
ZTF19aailpwl | 216.94340 | 29.51061 | AT2019brs | ATLAS19fyh, Gaia19axp | 19.3 | 0.3736 f |
ZTF20abgxlut | 257.76525 | 6.736331 | AT2020oio | Gaia20dvv, ATLAS20rmk | 20.4 | 0.247g |
ZTF20aadesap | 232.76545 | 53.40535 | AT2022fpx | ATLAS22kjn, Gaia22cwy, PS23bdt | 17.9 | 0.073h |
ZTF18acvvudh | 117.7084916 | 1.358392278 | AT2018lcp | 18.6 | - | |
ZTF22aaaeons | 231.827862 | -8.532609 | AT2022zg | Gaia22aft, ATLAS22ery | 18.5 | - |
Supplementary spectroscopic ANTs | ||||||
ZTF20acvfraq | 349.72403 | -10.58489 | AT2020adpi | ATLAS20bjzp, Gaia21aid | 19.7 | 0.26 i |
ZTF19aatubsj | 257.27857 | 26.85569 | AT2019fdr | ATLAS19lkd | 18 | 0.2666 j,h |
ZTF20aanxcpf | 15.16508 | 39.70842 | AT2021loi | ATLAS21qje | 18 | 0.083 k |
@@ 2.1.1 Nuclear transients
ANTs are, by definition, located in galaxy nuclei. To select nuclear transients we use the sherlock sky context software (Smith et al., 2020) running in Lasair. The ZTF search area is fully covered by the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS1) science consortium surveys (Chambers et al., 2016; Wainscoat et al., 2016): sherlock uses the Pan-STARRS catalogues and probabilistic classifications of unresolved point sources (Tachibana & Miller, 2018) to determine a host galaxy and the transient location within that host. We select all objects that are predicted as Nuclear Transients (i.e. transients consistent with a galactic nucleus) or AGN (transients consistent with a galactic nucleus that are listed in an AGN catalogue). Note that most transient searches exclude known AGN, but we allow them to avoid excluding bona-fide transient events associated with actively accreting SMBHs. We also include all objects predicted as SNe (transients not entirely consistent with a galaxy nucleus) with an angular separation of from the catalogued centre of a galaxy, based on the requirements used for the ZTF TDE samples (van Velzen et al., 2021; Yao et al., 2023a). We verify the nuclear origin of any transients in our sample by validating that their mean position lies within of the catalogued galaxy position, roughly equivalent to the ZTF positional uncertainty. Finally, we include “Orphan" transients, which have no associated host galaxy in deeper survey imaging (as was the case for AT2021lwx). Including transient that appear hostless prevents the systematic exclusion of events at high-redshift, extremely dusty, or uncatalogued dwarf galaxies.
@@ 2.2 Photometric selection of ANTs
Duration: ANTs are long-lived transients, and we require that the most recent detection in at least one filter must be at least 300 d after the first detection.
@@
Single, smooth decline: Any remaining AGN and variable stars should show stochastic variability on timescales shorter than the typical decline of an ANT , a property we denote ‘bumpiness’. We define the opposite, ‘smoothness’, as the lack of variability above the photometric noise on timescales shorter than the main lightcurve. Single transient events show power-law or exponential decays, which are linear in magnitude space. We fit the declines with a linear model and quantify the goodness-of-fit to the linear model with the Pearson parameter. We set an arbitrary threshold of for a decline to be accepted as smooth and linear. We vary this definition of ‘smoothness’ below and determine its effect on the number of high-quality candidates to be minimal.
The distribution of candidates in these parameter spaces is shown in Fig. 1. After all the nominal selection criteria are applied, we are left with 5855559 separate candidate light curves passed, but we note that ZTF22aadesap and ZTF22aafuzjv are two identifiers for the same light curve, leaving 58 distinct transients. light curves of nuclear transients and one orphan (ZTF20abrbeie=AT2021lwx), which we denote the ‘gold selection’. Of the ANTs published by Frederick et al. (2021) we recover ZTF19aailpwl=AT2019brs (the other three are too ‘bumpy’). To explore ANTs with different light curve shapes, we subsequently trial a set of different selection criteria, which are described in Appendix A. Briefly: we conduct a search for faster-declining events by relaxing the maximum duration to 180 d (the ‘faster’ sample); for less smooth events by relaxing the linear decline correlation coefficient to (the ‘bumpier’ sample); and the decline time to be longer than the rise time (the ‘platinum’ sample). The resulting samples and their ZTF light curve properties are available online666https://github.com/wisemanp/ANTs-Nest. For the 58 gold events, we obtain forced photometry for the full ZTF survey duration via the ZTF forced photometry service (Masci et al., 2019) for visual inspection.
@@ 2.2.1 Visual inspection
We visually inspect all 58 lightcurves and assign visual classifications of ‘Transient’, ‘AGN’, or ‘Unclear’ according to the flowchart in Fig 2. These classifications are predominantly based on pre-and post-peak variability. In particular, the forced photometry allows us to view any variability that occurred before the candidate passed the alert threshold in ZTF. We classify 16 events as Transients, six as Unclear, and 36 as AGN.
2.2.2 Spectroscopic history and follow-up
For the analysis in this paper, we require spectroscopy for two reasons: firstly it is needed to remove non-ambiguous transients that were selected photometrically (i.e. supernovae and TDEs), and secondly in order to estimate the absolute luminosity, energetics, and the elemental composition of ANTs. Of the 58 ‘gold’ light curves, 18 were reported as transients to the Transient Name Server (TNS)777https://www.wis-tns.org/, and a handful have been publicly spectroscopically classified. Note that ANT is not a spectroscopic classification. Here we cross-match the candidates we photometrically classified as transients with the public classifications. In the cases where the spectra are clearly supernovae, we discard the events from our sample. We do not discard events with AGN-like spectra if the light curves are classified as Transient (these are, by definition, ANTs). We also check transients labelled as Unclear – if they have AGN-like spectra we discard them as long-timescale AGN variability.
Of the 16 long-duration events we classified as transients via visual inspection, five are spectroscopically classified supernovae according to the public TNS classifications: ZTF19abpvbzf (SN IIn; Brennan et al. 2019), ZTF22aanwibf (SN II; Pellegrino et al. 2022). We also recovered the TDE ZTF21aaaokyp=AT2021axu (Hammerstein et al., 2021; Yao et al., 2023a) which we discard as it is not ambiguous.
The remaining events are detailed in Table 3. ZTF22aadesap=AT2022fpx (Perez-Fournon et al., 2022): The spectra of ZTF22aadesap=AT2022fpx do not resemble the main classes of TDE defined by Arcavi et al. (2014); van Velzen et al. (2021) and the object has recently been identified as an extreme coronal line emitter (Koljonen et al., 2024). The classification is thus ambiguous, so we include ZTF22aadesap=AT2022fpx as an ANT.
Seven of the remaining nine transients have been observed spectroscopically, four of which have public classifications while the other three are presented here for the first time. ZTF19aailpwl=AT2019brs is an ANT (Frederick et al., 2021), and ZTF20abrbeie=AT2021lwx (Wiseman et al., 2023; Subrayan et al., 2023) passes our selection by construction. ZTF21abxowzx=AT2021yzu and ZTF20abgxlut are spectroscopically classified as AGN at (Chu et al., 2021a) and respectively, but the light curves are clearly transient: in Section 3.1 we classify these as ANTs.
Further to public classifications, we obtained spectra for three that we photometrically and visually classified as Transients: ZTF18aczpgwm=AT2019kn, ZTF19aamrjar, and ZTF20abodaps=AT2020afep. These spectra are described in detail in Section 3.1. We also obtained spectra for three events which we labelled as Unclear: ZTF19aaozooc, ZTF20aaqtncr, ZTF20aauowyg. We discuss these events in Section 3.2. Finally, ZTF18acvvudh and ZTF22aaaeons have to our knowledge not been spectroscopically observed. They are therefore not considered for the remainder of this analysis, but we include them in Table 3 for completeness.